1/2x-5/3=3/4: 1/2x-5/3=3/4
Answer by jojo14344(883)

1/2x-5/3=3/4
(1/2)x-(5/3)=3/4
(1/2)x=(3/4)+(5/3)
(1/2)x=(9+20)/12 ---> (1/2)x=29/12
x=(29/12)/(1/2) ----> x=(29/12)(2/1)
x=4-5/6 ------------> final answer
Thank you,
Jojo